OpenStudy (anonymous):

solve the formula for area of a trapezoid. A = h(b1 + b2/2) for b1

OpenStudy (diyadiya):

\[A= h( \frac{b1+b2}{2})\]first multiply both sides by 2 can you do that?

OpenStudy (apoorvk):

You mean: \[A = h(\frac{b_1 + b_2} 2) \] where \(b_1~ and~b_2\) are length of parallel sides, and 'h' is the height right? Can you just rearrange the whole thing, to write it in terms of \(b_1\) solely?

OpenStudy (anonymous):

That is what I meant.

OpenStudy (apoorvk):

I mean, to write it *implicitly* for \(b_1\).

OpenStudy (diyadiya):

\[2 \times A = 2 \times h \left( \frac{b1+b2}{2} \right)\]\[2 A = \cancel{2} \times h \left( \frac{b1+b2}{\cancel{2}} \right)\]\[2 A = h \left( {b1+b2} \right)\] Can you continue from here?

OpenStudy (anonymous):

yeah thanks.

OpenStudy (anonymous):

I mean it's still kind of confusing..

OpenStudy (diyadiya):

Welcome!! But what will be the next step?

OpenStudy (anonymous):

well the answer is for b1

OpenStudy (anonymous):

I have no clue :(

OpenStudy (diyadiya):

lets go step by step though did you understand till this part?

OpenStudy (anonymous):

yeah pretty much

OpenStudy (diyadiya):

Sure?

OpenStudy (diyadiya):

Please do ask if you haven't got any specific step!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get it to this part.

OpenStudy (diyadiya):

ok so, 2A = h(b1 + b2) Now next step is to divide both sides by 'h' can you do that?

OpenStudy (anonymous):

is it b1 = a/h?

OpenStudy (diyadiya):

Nope, don't be in a hurry! do the step i told you to do :)

OpenStudy (anonymous):

okay okay :P

OpenStudy (diyadiya):

??

OpenStudy (anonymous):

yeah I can only do so much for an 8th grader lol

OpenStudy (anonymous):

Im in an advanced math and i'm horrible at it.

OpenStudy (diyadiya):

Its just division just divide both sides by h for example when you have c = bk and when you divide both sides by k you get c/k = bk/k c/k = b just like that!